# Nudged-Elastic Band Calculations of Polymorph Transitions and Solid-State Reactions in Molecular Crystals

### v1

@@ III Results and Discussion
Fig. 3 shows the obtained NEB results for several methods for the (left) to the (right) transition. The corresponding energy barriers (calculated from polymorph to the maximal energy) and the thermodynamic relative stability of the polymorphs , where a negative value indicates that polymorph is more stable, are listed in Table 1. The PBE+MBD/light results yield a small energy barrier of only 6.3 kJ/mol, which fits well with a transition close to room temperature. In terms of thermodynamics, both polymorphs have virtually the same stability, since we cannot properly distinguish stability differences of less than 1 kJ/mol. Please note, that the values reported here are relative static lattice energies (normalized per molecular unit), which do not include vibrational free energy effects. When performing tight PBE+MBD single points or tight PBE0+MBD estimates via multimer embeddingHoja, List, and Boese (2023) including up to trimer interactions (ME3), the relative energies change less than 1 kJ/mol.
@@
When analyzing the NEB results obtained with the semi-empirical density functional tight binding approach DFTB3+MBD, one can clearly see that there are significant differences compared to PBE+MBD. It can be seen that the energy barrier is with 4 kJ/mol much smaller and the energy difference between the two polymorphs is now larger than 3 kJ/mol. When we utilize the DFTB3+MBD structures and perform on top of them PBE+MBD/light single point calculations, we get similar results to the PBE+MBD NEB results, but the maximum is slightly shifted. This suggests that DFTB+MBD is not reliable in this context when it comes to energies, but the structures along the minimum energy path seem to have a better quality.
@@
Our SO3krates MLFF performed very well, the differences to the PBE+MBD NEB results and subsequent PBE+MBD single point calculations are less than 0.5 kJ/mol. When we analyze the differences between the SO3krates results and the PBE+MBD single points on the same structures along the minimum energy path, we obtain a mean absolute error (MAE) and mean error (ME) for the relative energies of only 0.2 and 0.1 kJ/mol, respectively. Performing the NEB calculations with 15 instead of the otherwise used 10 images, yields virtually identical results (within 0.1 kJ/mol). Hence, our MLFF yields very accurate results and by far outperformed DFTB3+MBD. Therefore, we have performed the full NEB calculations for the remaining two systems only with the generated MLFF.
@@
Since it is possible for this system to utilize a simple linear interpolation as starting point of NEB, we have compared the efficiency with our new interpolation method. Both methods yield identical results, but our new methodology converged slightly faster in 182 steps compared to 206 steps of the simple linear interpolation.
@@
After having discussed the energetics, we analyze the obtained transition-state geometries. The unit cell of the PBE+MBD transition state is visualized in Fig. 1. It can be seen that the orientation of the left two molecules is exactly in between both polymorphs. To assess the quality of our PBE+MBD transition-state geometry, we have calculated the phonon frequencies and find two imaginary frequencies at the -point: -73.6 cm-1, which corresponds to the movement along the NEB pathway, and another mode at -14.5 cm-1. At a true transition state, we should only observe the first imaginary mode. Therefore, we have slightly displaced the structure along the second imaginary mode, followed by a subsequent phonon calculation. In this way we were able to fully remove this second imaginary mode and have obtained a true transition state. The differences between the two structures are essentially negligible, with a lattice energy difference of 0.014 kJ/mol and a RMSD20 value of 0.038 Å, which is the RMSD value of a cluster consisting of 20 molecules. When we compare the obtained transition states of DFTB3+MBD and SO3krates with the PBE+MBD results, we obtain RMSD20 values of 0.308 Å and 0.193 Å, respectively. Also in this case, the MLFF outperforms DFTB3+MBD and yields a structure very close to the PBE+MBD result.
@@ III.2 Theophylline Polymorph Transition
Next, we discuss the polymorphic transition from theophylline polymorph I Khamar et al. (2011), which is stable at higher temperatures, to the room-temperature stable polymorph II Fucke et al. (2012b). This transition (I to II) does not occur upon cooling due to significant structural differences.Fucke et al. (2012a) However, the reverse transition of polymorph II into I is known to occur upon heating. Since we can expect here a lot of atomic motion, this system is ideal for testing our new interpolation method. When dividing every molecule into three fragments (see Fig. 2), our new interpolation method automatically creates initial reaction path structures without any atomic collisions.
The results of the minimum-energy path obtained with the SO3krates MLFF together with single point calculations on top of the SO3krates structures are shown in Fig. 4. In both cases, polymorph II is, as expected, thermodynamically more stable by about 5 kJ/mol. In both cases we obtain a substantial energy barrier, amounting to 46.7 kJ/mol and 49.7 kJ/mol for SO3krates and the PBE+MBD single points, respectively. SO3krates provides again very accurate results, underestimating the PBE+MBD energy barrier by only 6 % and with a MAE and ME across the minimum energy path of 0.9 and 0.8 kJ/mol, respectively. The obtained transition state is shown in Fig. 1 and one can see that this process involves substantial motion of all molecules, given the large changes in unit-cell parameters.
@@ III.3 Solid-State Diels-Alder Reaction
Finally, we investigate if our approach can also be applied to solid-state reactions. As example we utilize a single-crystal to single-crystal Diels-Alder reaction of a co-crystal composed of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene.Khorasani et al. (2015) Experimentally, it took 84 days to achieve a 97 % conversion.
When dividing both molecules building up the co-crystal into several fragments (see Fig. 2), our interpolation yielded also for this reaction automatically collision-free initial path structures. Fig. 5 shows again the results of the minimum-energy path obtained with the SO3krates MLFF together with single point calculations on top of the SO3krates structures. In both cases, the product of the reaction is about 5.5 kJ/mol more stable than the co-crystal formed by the reactants. The energy barrier is for both methods very similar amounting for SO3krates and PBE+MBD to 27.7 and 28.6 kJ/mol, respectively. The deviation of the MLFF energy barrier compared to the DFT reference is less than 1 kJ/mol and statistically over the whole minimum energy path the MAE and ME amount to only 0.5 and -0.1 kJ/mol, respectively. The corresponding transition state is visualized in Fig. 1. Before the actual ring closure can happen, several propenyl groups need to re-arrange themselves, which apparently is the main source of the energy barrier in this case.
@@ IV Conclusion
We have introduced a novel interpolation method for generating an initial pathway for subsequent nudged-elastic band (NEB) calculations suitable for periodic structures like molecular crystals. This methodology aims to automatically avoid any atomic collisions, which otherwise regularly happen in periodic systems when using simple linear interpolations. We demonstrated the versatility of this approach by successfully applying it to a range of divers systems, including two polymorph transitions — one with a small and one with a rather large energy barrier — as well as a single-crystal Diels-Alder reaction. Our approach consistently provided collision-free initial pathways, avoiding the need for cumbersome manual corrections. This automated pathway generation greatly simplifies the initial setup for NEB calculations, making it a highly practical tool for tackling even complex periodic systems.
Furthermore, we have utilized the SO3krates architecture to train an individual machine-learned force field (MLFF) for every studied system. By training the MLFFs on a diverse set of reference data obtained from PBE+MBD calculations, we were able to reproduce the energy barriers and structural characteristics of the minimum energy pathways with high accuracy. The MAEs between the MLFF energies along the reaction paths compared to subsequent PBE+MBD single-point energies remained consistently below 1 kJ/mol for all studied systems. However, we note that the reference data focused heavily on the PES region required for the studied transition and the resulting MLFFs will likely be much less accurate when simulating for instance other polymorphs.
Finally, many polymorph transition and reaction paths can probably not be described by utilizing unit cells but would require larger supercells to allow for the needed flexibility. In such cases, computationally efficient methods like MLFFs are key for being able to afford the necessary calculations. Our results highlight that with properly trained MLFFs, it is possible to achieve both computational efficiency and accuracy, making them a very promising tool for exploring complex solid-state phenomena.
@@ Acknowledgements.
The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC).References

### v2

@@ III Results and Discussion
Fig. 3 shows the obtained NEB results for several methods for the transition of the polymorph (left) to the polymorph (right). The corresponding energy barriers (calculated from polymorph to the maximal energy) and the thermodynamic relative stability of the polymorphs , where a negative value indicates that polymorph is more stable, are listed in Table 1. The small unit cell size and the low complexity of this polymorph transition allow us to perform all necessary calculations fully at the DFT level. Hence, we utilize 1,4-diiodobenzene to properly evaluate the performance of our SO3krates MLFF for energies and structures before applying the MLFF models to more complex systems. The PBE+MBD/light results yield a small energy barrier of only 6.3 kJ mol-1, which fits well with a transition temperature close to room temperature. In terms of thermodynamics, both polymorphs have virtually the same stability ( kJ mol-1), since we cannot reliably distinguish between such small stability differences with DFT calculations. Please note that the values reported here are relative static lattice energies (normalized per molecular unit), which do not include vibrational free energy effects. When performing tight PBE+MBD single points or tight PBE0+MBD estimates via multimer embeddingHoja, List, and Boese (2023) including up to trimer interactions (ME3), the relative energies change less than 1 kJ mol-1.
@@
When analyzing the NEB results obtained with the semi-empirical density functional tight binding approach (DFTB3+MBD), one can clearly see that there are significant differences compared to PBE+MBD. The energy barrier is with 4 kJ mol-1 much smaller and the energy difference between the two polymorphs is now much larger, amounting to more than 3 kJ mol-1. When we utilize the DFTB3+MBD structures and perform on top of them PBE+MBD/light single-point calculations, we get similar results to the PBE+MBD NEB results, but the maximum is slightly shifted. This suggests that DFTB3+MBD is not reliable in this context when it comes to energies, but the structures along the minimum energy path seem to have a better quality.
@@
Next, we evaluate the performance of our SO3krates MLFF. When we compare the and values obtained with SO3krates with those of the fully optimized PBE+MBD/light calculations, both values differ by less than 0.3 kJ mol-1. Also, the maximum along the minimum-energy path is located at the same image in both cases and the energy difference of the two methods for the individual points along the path remains consistently below 0.6 kJ mol-1. Hence, SO3krates is able to very accurately reproduce the PBE+MBD energies of this transition. Additionally, we have performed PBE+MBD single point calculations on top of the SO3krates structures to evaluate the accuracy of the energies without the influence of geometry effects. In this case, all differences between the SO3krates results and the PBE+MBD single points (on the same structures) along the minimum energy path are less than 0.5 kJ mol-1 and we obtain a mean absolute error (MAE) and a mean error (ME) for the relative energies of only 0.2 and -0.1 kJ mol-1, respectively. Performing the NEB calculations with 15 instead of the otherwise used 10 images yields virtually identical results (within 0.1 kJ mol-1). Hence, our SO3krates MLFF yields very accurate results and significantly outperforms DFTB3+MBD. Therefore, we have performed the full NEB calculations for the remaining two systems only with the generated MLFF.
However, the so far discussed accuracies of the obtained SO3krates MLFF were only based on a few energy differences. Therefore, Table 2 lists the accuracy of the MLFF for a large test set containing 7125 1,4-diiodobenzene structures, which were not used for training or validation. It can be seen that the obtained energies and forces are highly accurate. The MAE of the energies amounts to only 0.13 kJ mol-1 per molecular unit or 0.01 kJ mol-1 when expressed per atom, and the MAE of the forces amounts to only 1.1 kJ mol-1 Å-1, which corresponds to a coefficient of determination of 0.9998. To put the force error into perspective, it corresponds to 0.011 eV Å-1, which is only about twice as large as our (quite strict) convergence setting used for the DFT optimizations (0.005 eV Å-1).
| Energies | Forces | |||
| System | MAE | R2 | MAE | R2 |
| 1,4-Diiodobenzene | 0.13 | 0.9998 | 1.10 | 0.9998 |
| Theophylline | 0.26 | 0.9999 | 1.68 | 0.9994 |
| Diels-Alder Reaction | 0.75 | 0.9996 | 1.92 | 0.9993 |
@@
Since it is possible for this system to utilize a simple linear interpolation as starting point of NEB, we have compared the efficiency with our new interpolation method. Both methods yield identical results, but our new methodology converged slightly faster, needing 12 % fewer steps than the simple linear interpolation.
@@
After having discussed the energetics, we analyze the obtained transition-state geometries. The unit cell of the PBE+MBD transition state is visualized in Fig. 1. It can be seen that the orientation of the left two molecules is exactly in between both polymorphs. To assess the quality of our PBE+MBD transition-state geometry, we have calculated the phonon frequencies and find two imaginary frequencies at the -point: -73.6 cm-1, which corresponds to the movement along the NEB pathway, and another mode at -14.6 cm-1. At a true transition state, we should only observe the first imaginary mode. Therefore, we have slightly displaced the structure along the second imaginary mode, followed by a subsequent phonon calculation. In this way, we were able to fully remove this second imaginary mode and have obtained a true transition state for this polymorph transition. The differences between the two structures are essentially negligible, with a lattice energy difference of 0.01 kJ mol-1 and a RMSD20 value of 0.038 Å, which is the root mean-square deviation (RMSD) of a cluster consisting of 20 molecules.
In order to illustrate the effect of vibrational free energies on and , we have additionally performed PBE+MBD/light phonon calculations for the optimized structures of polymorphs and . At room temperature, the resulting free energy difference in polymorph stability amounts to -0.21 kJ mol-1. This is only a very small shift by 0.01 kJ mol-1 compared to the static lattice energies, but we note that for other systems the inclusion of vibrational free energies can often be crucial for determining the correct thermodynamic stabilitiesHoja et al. (2019). The calculated free energy barrier amounts at room temperature to 5.4 kJ mol-1, which is about 1 kJ mol-1 lower than the corresponding static lattice energy result.
When we compare the obtained transition states of DFTB3+MBD and SO3krates with the PBE+MBD results, we obtain RMSD20 values of 0.308 Å and 0.199 Å, respectively. Also in this case, the SO3krates MLFF outperforms DFTB3+MBD and yields a structure very close to the PBE+MBD result. A detailed analysis of the accuracy of the MLFF for phonon calculations and transition state optimizations will follow in a subsequent publication.
@@ III.2 Theophylline Polymorph Transition
Next, we discuss the polymorphic transition from theophylline polymorph I Khamar et al. (2011), which is stable at higher temperatures, to the room-temperature stable polymorph II Fucke et al. (2012b). This transition (I to II) does not occur upon cooling due to significant structural differences.Fucke et al. (2012a) However, the reverse transition of polymorph II into I is known to occur upon heating. Since we can expect here a lot of atomic motion, this system is ideal for testing our new interpolation method. When dividing every molecule into three fragments (see Fig. 2), our new interpolation method automatically creates initial reaction path structures without any atomic collisions, while this is not possible with a standard linear interpolation.
The results of the minimum-energy path obtained with the SO3krates MLFF together with single point calculations on top of the SO3krates structures are shown in Fig. 4. In both cases, polymorph II is, as expected, thermodynamically more stable by about 5 kJ mol-1 and we obtain a substantial energy barrier, amounting to 46.4 kJ mol-1 and 47.7 kJ mol-1 for SO3krates and the PBE+MBD single points, respectively. The SO3krates MLFF provides again very accurate results, underestimating the PBE+MBD energy barrier by only 3 % and yielding MAE and ME values across the minimum energy path of 0.3 and -0.2 kJ mol-1, respectively. The obtained transition state is shown in Fig. 1 and one can see that this process involves substantial motion of all molecules, given the large changes in unit-cell parameters.
The MAE of our theophylline SO3krates MLFF across a test set containing 9276 structures amounts for energies to 0.26 kJ mol-1 when normalized per molecule (see Table 2) and again 0.01 kJ mol-1 when normalized per atom. The MAE of the forces is with 1.68 kJ mol-1 Å-1 only slightly higher than that for 1,4-diiodobenzene.
@@ III.3 Solid-State Diels-Alder Reaction
Finally, we investigate if our approach can also be applied to solid-state reactions. As an example we utilize a single-crystal to single-crystal Diels-Alder reaction of a co-crystal composed of bis(N-allylimino)-1,4-dithiin and 9-bromoanthracene.Khorasani et al. (2015) Experimentally, it took 84 days to achieve a 97 % conversion.
When dividing both molecules building up the co-crystal into several fragments (see Fig. 2), our interpolation yielded also for this reaction automatically collision-free initial path structures. Fig. 5 shows the results of the minimum-energy path obtained with the SO3krates MLFF together with subsequent single point calculations on top of the SO3krates structures. In both cases, the product of the reaction is about 6.0 kJ mol-1 more stable than the co-crystal formed by the reactants. The energy barrier is for SO3krates and PBE+MBD virtually identical and amounts to 27.5 kJ mol-1 in both cases. When comparing the differences between the two methods for the individual images along the obtained minimum energy path, we find a maximum deviation of 1.5 kJ mol-1 and the MAE and ME amount to only 0.4 and -0.2 kJ mol-1, respectively. Hence, our obtained SO3krates MLFF is also able to describe the energetics along the minimum energy path of a solid-state reaction with high accuracy.
The corresponding transition state is visualized in Fig. 1. Before the actual ring closure can happen, several propenyl groups need to rearrange themselves, which apparently is the main source of the energy barrier in this case.
After having discussed the specific minimum energy path for the reaction, we illustrate the accuracy of the SO3krates model for a large test set containing 8165 structures (see Table 2). This test set includes co-crystal structures of bis(N-allylimino)-1,4-dithiin/9-bromoanthracene as well as structures containing the respective Diels-Alder product. Therefore, the molecular unit we have used to normalize the energies corresponds to one product molecule or one co-crystal unit. The resulting MAE amounts to 0.75 kJ mol-1, which is very accurate, especially given the already quite complicated system. While the error appears larger than that for the other two systems in this notation (normalized per molecular unit), it is virtually identical to that of the other systems when normalized per atom (0.01 kJ mol-1). The forces have with 1.92 kJ mol-1 Å-1 a slightly larger MAE than the two previous systems, but are with a coefficient of determination of 0.9993 still very accurate.
@@ IV Conclusion
We have introduced an interpolation method for generating an initial pathway for subsequent nudged-elastic band (NEB) calculations suitable for periodic structures like molecular crystals. This methodology aims to automatically avoid any atomic collisions, which otherwise regularly happen in periodic systems when using simple linear interpolations. We demonstrated the versatility of this approach by successfully applying it to a range of diverse systems, including two polymorph transitions — one with a small and one with a rather large energy barrier — as well as a single-crystal Diels-Alder reaction. Our approach consistently provided collision-free initial pathways, avoiding the need for cumbersome manual corrections. This automated pathway generation greatly simplifies the initial setup for NEB calculations, making it a highly practical tool for tackling even complex periodic systems.
Furthermore, we have utilized the SO3krates architecture to train an individual machine-learned force field (MLFF) for every studied system. By training the MLFFs on a diverse set of reference data obtained from PBE+MBD calculations for the specific systems, we are able to reproduce the energy barriers and structural characteristics of the minimum energy pathways with high accuracy. The MAEs between the MLFF energies along the reaction paths compared to subsequent PBE+MBD single-point energies are for all three systems smaller than 0.4 kJ mol-1 and the maximal error for individual images amounts to less than 1.5 kJ mol-1. In terms of our respective test sets containing in all cases more than 7000 structures calculated with PBE+MBD, all our MAEs for energies normalized per molecular unit amount to less than 0.8 kJ mol-1 or when expressed as MAE per atom consistently yield an error of 0.01 kJ mol-1. In terms of the forces, all three MAEs are below 2 kJ mol-1 Å-1. However, we note that the reference data focused heavily on the PES region required for the studied transition and the resulting MLFFs will likely be less accurate when simulating for instance other polymorphs.
Finally, many polymorph transition and reaction paths can probably not be described by utilizing unit cells but would require larger supercells to allow for the needed flexibility. In such cases, computationally efficient methods like MLFFs are key for being able to afford the necessary calculations. Our results highlight that with properly trained MLFFs, it is possible to achieve both computational efficiency and accuracy, making them a very promising tool for exploring complex solid-state phenomena. Therefore, the next steps will be to apply this methodology to NEB calculations involving larger supercells, to utilize these MLFFs for explicit molecular dynamics simulations, and to test their applicability for vibrational free energy calculations of molecular crystals.
@@ Acknowledgements.
The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC) and we thank Prof. A. Daniel Boese for providing resources at the University of Graz.References
@@
- Bauer et al. (2001) J. Bauer, S. Spanton, R. Henry, J. Quick, W. Dziki, W. Porter, and J. Morris, Pharm. Res. 18, 859–866 (2001).
- Bučar, Lancaster, and Bernstein (2015) D. Bučar, R. W. Lancaster, and J. Bernstein, Angew. Chem. Int. Ed. 54, 6972–6993 (2015).
- Price (2014) S. L. Price, Chem. Soc. Rev. 43, 2098–2111 (2014).
- Reilly et al. (2016) A. M. Reilly, R. I. Cooper, C. S. Adjiman, S. Bhattacharya, A. D. Boese, J. G. Brandenburg, P. J. Bygrave, R. Bylsma, J. E. Campbell, R. Car, D. H. Case, R. Chadha, J. C. Cole, K. Cosburn, H. M. Cuppen, F. Curtis, G. M. Day, R. A. DiStasio Jr, A. Dzyabchenko, B. P. van Eijck, D. M. Elking, J. A. van den Ende, J. C. Facelli, M. B. Ferraro, L. Fusti-Molnar, C.-A. Gatsiou, T. S. Gee, R. de Gelder, L. M. Ghiringhelli, H. Goto, S. Grimme, R. Guo, D. W. M. Hofmann, J. Hoja, R. K. Hylton, L. Iuzzolino, W. Jankiewicz, D. T. de Jong, J. Kendrick, N. J. J. de Klerk, H.-Y. Ko, L. N. Kuleshova, X. Li, S. Lohani, F. J. J. Leusen, A. M. Lund, J. Lv, Y. Ma, N. Marom, A. E. Masunov, P. McCabe, D. P. McMahon, H. Meekes, M. P. Metz, A. J. Misquitta, S. Mohamed, B. Monserrat, R. J. Needs, M. A. Neumann, J. Nyman, S. Obata, H. Oberhofer, A. R. Oganov, A. M. Orendt, G. I. Pagola, C. C. Pantelides, C. J. Pickard, R. Podeszwa, L. S. Price, S. L. Price, A. Pulido, M. G. Read, K. Reuter, E. Schneider, C. Schober, G. P. Shields, P. Singh, I. J. Sugden, K. Szalewicz, C. R. Taylor, A. Tkatchenko, M. E. Tuckerman, F. Vacarro, M. Vasileiadis, A. Vazquez-Mayagoitia, L. Vogt, Y. Wang, R. E. Watson, G. A. de Wijs, J. Yang, Q. Zhu, and C. R. Groom, Acta Crystallogr. Sect. B: Struct. Sci. 72, 439–459 (2016).
- Hunnisett et al. (2024a) L. M. Hunnisett, J. Nyman, N. Francia, N. S. Abraham, C. S. Adjiman, S. Aitipamula, T. Alkhidir, M. Almehairbi, A. Anelli, D. M. Anstine, J. E. Anthony, J. E. Arnold, F. Bahrami, M. A. Bellucci, R. M. Bhardwaj, I. Bier, J. A. Bis, A. D. Boese, D. H. Bowskill, J. Bramley, J. G. Brandenburg, D. E. Braun, P. W. V. Butler, J. Cadden, S. Carino, E. J. Chan, C. Chang, B. Cheng, S. M. Clarke, S. J. Coles, R. I. Cooper, R. Couch, R. Cuadrado, T. Darden, G. M. Day, H. Dietrich, Y. Ding, A. DiPasquale, B. Dhokale, B. P. van Eijck, M. R. J. Elsegood, D. Firaha, W. Fu, K. Fukuzawa, J. Glover, H. Goto, C. Greenwell, R. Guo, J. Harter, J. Helfferich, D. W. M. Hofmann, J. Hoja, J. Hone, R. Hong, G. Hutchison, Y. Ikabata, O. Isayev, O. Ishaque, V. Jain, Y. Jin, A. Jing, E. R. Johnson, I. Jones, K. V. J. Jose, E. A. Kabova, A. Keates, P. F. Kelly, D. Khakimov, S. Konstantinopoulos, L. N. Kuleshova, H. Li, X. Lin, A. List, C. Liu, Y. M. Liu, Z. Liu, Z.-P. Liu, J. W. Lubach, N. Marom, A. A. Maryewski, H. Matsui, A. Mattei, R. A. Mayo, J. W. Melkumov, S. Mohamed, Z. Momenzadeh Abardeh, H. S. Muddana, N. Nakayama, K. S. Nayal, M. A. Neumann, R. Nikhar, S. Obata, D. O’Connor, A. R. Oganov, K. Okuwaki, A. Otero-de-la Roza, C. C. Pantelides, S. Parkin, C. J. Pickard, L. Pilia, T. Pivina, R. Podeszwa, A. J. A. Price, L. S. Price, S. L. Price, M. R. Probert, A. Pulido, G. R. Ramteke, A. U. Rehman, S. M. Reutzel-Edens, J. Rogal, M. J. Ross, A. F. Rumson, G. Sadiq, Z. M. Saeed, A. Salimi, M. Salvalaglio, L. Sanders de Almada, K. Sasikumar, S. Sekharan, C. Shang, K. Shankland, K. Shinohara, B. Shi, X. Shi, A. G. Skillman, H. Song, N. Strasser, J. van de Streek, I. J. Sugden, G. Sun, K. Szalewicz, B. I. Tan, L. Tan, F. Tarczynski, C. R. Taylor, A. Tkatchenko, R. Tom, M. E. Tuckerman, Y. Utsumi, L. Vogt-Maranto, J. Weatherston, L. J. Wilkinson, R. D. Willacy, L. Wojtas, G. R. Woollam, Z. Yang, E. Yonemochi, X. Yue, Q. Zeng, Y. Zhang, T. Zhou, Y. Zhou, R. Zubatyuk, and J. C. Cole, Acta Crystallogr. Sect. B: Struct. Sci. 80, 517–547 (2024a).
- Hunnisett et al. (2024b) L. M. Hunnisett, N. Francia, J. Nyman, N. S. Abraham, S. Aitipamula, T. Alkhidir, M. Almehairbi, A. Anelli, D. M. Anstine, J. E. Anthony, J. E. Arnold, F. Bahrami, M. A. Bellucci, G. J. O. Beran, R. M. Bhardwaj, R. Bianco, J. A. Bis, A. D. Boese, J. Bramley, D. E. Braun, P. W. V. Butler, J. Cadden, S. Carino, C. Červinka, E. J. Chan, C. Chang, S. M. Clarke, S. J. Coles, C. J. Cook, R. I. Cooper, T. Darden, G. M. Day, W. Deng, H. Dietrich, A. DiPasquale, B. Dhokale, B. P. van Eijck, M. R. J. Elsegood, D. Firaha, W. Fu, K. Fukuzawa, N. Galanakis, H. Goto, C. Greenwell, R. Guo, J. Harter, J. Helfferich, J. Hoja, J. Hone, R. Hong, M. Hušák, Y. Ikabata, O. Isayev, O. Ishaque, V. Jain, Y. Jin, A. Jing, E. R. Johnson, I. Jones, K. V. J. Jose, E. A. Kabova, A. Keates, P. F. Kelly, J. Klimeš, V. Kostková, H. Li, X. Lin, A. List, C. Liu, Y. M. Liu, Z. Liu, I. Lončarić, J. W. Lubach, J. Ludík, N. Marom, H. Matsui, A. Mattei, R. A. Mayo, J. W. Melkumov, B. Mladineo, S. Mohamed, Z. Momenzadeh Abardeh, H. S. Muddana, N. Nakayama, K. S. Nayal, M. A. Neumann, R. Nikhar, S. Obata, D. O’Connor, A. R. Oganov, K. Okuwaki, A. Otero-de-la Roza, S. Parkin, A. Parunov, R. Podeszwa, A. J. A. Price, L. S. Price, S. L. Price, M. R. Probert, A. Pulido, G. R. Ramteke, A. U. Rehman, S. M. Reutzel-Edens, J. Rogal, M. J. Ross, A. F. Rumson, G. Sadiq, Z. M. Saeed, A. Salimi, K. Sasikumar, S. Sekharan, K. Shankland, B. Shi, X. Shi, K. Shinohara, A. G. Skillman, H. Song, N. Strasser, J. van de Streek, I. J. Sugden, G. Sun, K. Szalewicz, L. Tan, K. Tang, F. Tarczynski, C. R. Taylor, A. Tkatchenko, R. Tom, P. Touš, M. E. Tuckerman, P. A. Unzueta, Y. Utsumi, L. Vogt-Maranto, J. Weatherston, L. J. Wilkinson, R. D. Willacy, L. Wojtas, G. R. Woollam, Y. Yang, Z. Yang, E. Yonemochi, X. Yue, Q. Zeng, T. Zhou, Y. Zhou, R. Zubatyuk, and J. C. Cole, Acta Crystallogr. Sect. B: Struct. Sci. 80, 548–574 (2024b).
- Price (2013) S. L. Price, Acta Crystallogr. Sect. B: Struct. Sci. 69, 313–328 (2013).
- Price (2022) S. L. Price, Faraday Discuss. 235, 569–581 (2022).
- Beran (2023) G. J. O. Beran, Chem. Sci. 14, 13290–13312 (2023).
@@
- Francia et al. (2020) N. F. Francia, L. S. Price, J. Nyman, S. L. Price, and M. Salvalaglio, Cryst. Growth Des. 20, 6847–6862 (2020).
- Yang and Day (2022) S. Yang and G. M. Day, Commun. Chem. 5, 86 (2022).
- Butler and Day (2023) P. W. V. Butler and G. M. Day, Proc. Natl. Acad. Sci. 120, e2300516120 (2023).
@@
- Chmiela et al. (2017) S. Chmiela, A. Tkatchenko, H. E. Sauceda, I. Poltavsky, K. T. Schütt, and K.-R. Müller, Sci. Adv. 3, e1603015 (2017).
@@
- Chmiela et al. (2018) S. Chmiela, H. E. Sauceda, K.-R. Müller, and A. Tkatchenko, Nat. Commun. 9, 3887 (2018).
@@
- Poltavsky and Tkatchenko (2021) I. Poltavsky and A. Tkatchenko, J. Phys. Chem. Lett. 12, 6551–6564 (2021).
@@
- Sauceda et al. (2022) H. E. Sauceda, L. E. Gálvez-González, S. Chmiela, L. O. Paz-Borbón, K.-R. Müller, and A. Tkatchenko, Nat. Commun. 13, 3733 (2022).
@@
- Kovács et al. (2025) D. P. Kovács, J. H. Moore, N. J. Browning, I. Batatia, J. T. Horton, Y. Pu, V. Kapil, W. C. Witt, I.-B. Magdău, D. J. Cole, and G. Csányi, J. Am. Chem. Soc. 147, 17598–17611 (2025).
- Raimbault et al. (2019) N. Raimbault, A. Grisafi, M. Ceriotti, and M. Rossi, New J. Phys. 21, 105001 (2019).
- Krynski and Rossi (2021) M. Krynski and M. Rossi, Npj Comput. Mater. 7, 169 (2021).
- Lewis, Lazzaroni, and Rossi (2023) A. M. Lewis, P. Lazzaroni, and M. Rossi, J. Chem. Phys 159, 014103 (2023).
- Žugec, Geilhufe, and Lončarić (2024) I. Žugec, R. M. Geilhufe, and I. Lončarić, J. Chem. Phys. 160, 154106 (2024).
- Mladineo and Lončarić (2024) B. Mladineo and I. Lončarić, Cryst. Growth Des. 24, 8167–8173 (2024).
- Zhao et al. (2025) C. Zhao, H. Liu, D.-H. Qu, A. I. Cooper, and L. Chen, Chem. Sci. (2025), 10.1039/d4sc06467d.
- Nickerson and Johnson (2025) C. J. Nickerson and E. R. Johnson, Phys. Chem. Chem. Phys. 27, 11930–11940 (2025).
@@
- Gaus, Goez, and Elstner (2012) M. Gaus, A. Goez, and M. Elstner, J. Chem. Theory Comput. 9, 338–354 (2012).
- Gaus et al. (2014) M. Gaus, X. Lu, M. Elstner, and Q. Cui, J. Chem. Theory Comput. 10, 1518–1537 (2014).
- Kubillus et al. (2014) M. Kubillus, T. Kubař, M. Gaus, J. Řezáč, and M. Elstner, J. Chem. Theory Comput. 11, 332–342 (2014).